Find the sum of the sequence algebraically.
5k+3 is the common difference
k=1 is the starting term
Number of terms is 5

Answer by stanbon(75887)

Find the sum of the sequence algebraically.
5k+3 is the common difference
k=1 is the starting term
Number of terms is 5
------
a(5) = 1 + 4(5k+3) = 20k+13
-----
S(5) = (5/2)(1 + 20k+13)
---
S(5) = (5/2)(20k+14)
---
S(5) = 5(10k+7)
--
S(5) = 50k+35
==================
Cheers,
Stan H.
==================